OpenStudy (anonymous):

find an equation of the line that passes through(0,5) and is parallel to y=1/2x+4 equation of the line is..

OpenStudy (mathteacher1729):

You're looking for a line with slope 1/2 (why is the slope 1/2 ?) and that goes through the point (0,5). If you use the point-slope form, you can just plug in these numbers and you'll be done. :)

OpenStudy (anonymous):

So the slope of the lines would be the same (1/2) and so you would then use point-slope form: y-5=1/2(x-0) Once that is simplified ( y - 5 = 1/2x + 0 ) you convert it to slope-intercept form and you get y = 1/2x +5 That should be your answer.